In a charming 19th-century hotel, an old-style elevator is connected to a counterweight by a cable that passes over a rotating disk 1.87m

1.87m in diameter. The elevator is raised and lowered by turning the disk, and the cable does not slip on the rim of the disk but turns with it. Through how many revolutions has the disk turned when it has raised the elevator 3.36m

3.36m between floors?


1
Expert's answer
2022-04-17T17:06:05-0400

The length of the disk's circumference is:



"l = 2\pi R"

where "R = 1.87m\/2 = 0.935m". In order to make one revolution, it must pass the rope of length "l" through itself. Thus, if the elevator was raised by "L = 3.36m", the number of turn required is:



"N = \dfrac{L}{l} = \dfrac{L}{2\pi R}\\
N = \dfrac{3.36m}{2\pi \cdot 0.935m} \approx 0.572"

Answer. 0.572 revolutions.